# On Singular Integrals and Quantitative Rectifiability in Parabolic Space and the Heisenberg Group

### v1

@@ 5. Vampiric measures
In light of Theorem 4.3, the task of proving Theorems 3.1 and 3.2 have been reduced to proving Carleson packing conditions on the collection of cubes $E\in\mathbb{D}(\mu)$ with noticeable $\alpha_{\mu,k}(E)$ coefficient. These theorems are provided by studying vampiric measures. The majority of the analysis takes place in the parabolic setting. The situation with the Heisenberg group will then be reduced to the parabolic case by appealing to the analysis in Orponen [Orp]. It is important to note that we are not only interested in the support of a vampiric measure but also its distribution.
We will call a Borel measure $\mu$ on $\mathbb{R}^{n+1}$ vampiric [^8] These measures correspond to the local symmetry condition introduced by David and Semmes [DS1], and so it might be natural to refer to them as symmetric measures. However, Mattila has already introduced a (wider) class of measures in the context of understanding the geometric consequences of the existence of principal value singular integrals called symmetric measures, so we decided not to use this terminology. if
@@ 5.1. The parabolic case
For every $\lambda>0$ there exist $M_{1}>0$ and $M_{2}>0$ such that the following holds:
Suppose that a vampiric measure $\mu$ is $(k+2)$-ADR, $\mathbf{0}\in\operatorname{supp}(\mu)$ and
(5.2) $\alpha_{\mu,k}(B(\mathbf{0},r))>0.$
@@
(5.3) $\alpha_{\mu,k+1}(B(\mathbf{0},M_{1}r))<\lambda$
@@
(5.4) $\alpha_{\mu,k+2}(B(\mathbf{0},M_{2}r))<\lambda.$
The proof of this proposition will require a fair amount of preparatory work. Recall the definitions of cylinders $C_{r}(\mathbf{X})$ and spatial cubes $\widetilde{Q}(X,r)$ from Section 2.1.
@@
Suppose $(X,t),(Y,s)\in\operatorname{supp}(\mu)$. With Z=YXZ=Y-X, the points $(X+2lZ,t)$, $k\in\mathbb{Z}$, are contained in $\operatorname{supp}(\mu)$. Moreover, for all $0<r<\infty$ we have
$\mu\big(C_{r}(X+2l(Y-X),t)\big)=\mu\big(C_{r}(X,t)\big).$
@@
With a view to arguing by induction, consider the following statement: Suppose that $k\in\{0,1,2,\dots\,\}$ and that $(X+2\ell(Y-X),t)\in\operatorname{supp}(\mu)$ and $(Y+2\ell(Y-X),s)\in\operatorname{supp}(\mu)$ for every $\ell\in\mathbb{Z}$ with $|\ell|\leq k$, and $\mu\big(C_{r}(X+2l(Y-X),t)\big)=\mu\big(C_{r}(X,t)\big)$ for $\ell\in\mathbb{Z}$ with $|\ell|\leq k$.
This statement holds by assumption for $k=0$. Suppose now it holds for some fixed $k$, then
$C_{r}(X+2(k+1)(Y-X),t)=(Y+2k(Y-X),s)+C_{r}((Y-X),t-s),$
so equation (5.5) with $(X,t)$ replaced by $(Y+2k(Y-X),s)\in\operatorname{supp}(\mu)$, $\xi=Y-X$, and $\tau=t-s$, yields
$\mu\big(C_{r}(X+2(k+1)(Y-X),t)\big)$ $=\mu\big((Y+2k(Y-X),s)+C_{r}((X-Y),t-s)\big)$
$=\mu\big(C_{r}(X+2k(Y-X),t)\big)>0.$
@@
$C_{r}(X+2(k+1)(Y-X),t)=(X,t)+C_{r}(2(k+1)(Y-X),0),$
so employing (5.5), we deduce that $\mu(C_{r}(X+2(k+1)(Y-X),t))=\mu(C_{r}(X-2(k+1)(Y-X),t))>0.$ Interchanging the roles of $(X,t)$ and $(Y,s)$ in the above argument now yields that $(Y\pm 2(k+1)(Y-X),s)\in\operatorname{supp}(\mu)$. This completes the induction step, and proves the lemma. ∎
@@
Suppose $(X,t),(Y,t)\in\operatorname{supp}(\mu)$. With Z=YXZ=Y-X, the points $(X+lZ,t)$, $l\in\mathbb{Z}$, are contained in $\operatorname{supp}(\mu)$.
@@
Suppose $(X,t),(Y,s)\in\operatorname{supp}(\mu)$, $r>0$, and $I$ is an open interval in $\mathbb{R}$. With $\widetilde{Q}_{r}(Z):=\{Z^{\prime}\in\mathbb{R}^{n}:\|Z-Z^{\prime}\|_{\infty}<r\}$, for each $l\in\mathbb{Z}$ we have
$\mu(\widetilde{Q}_{r}(X+2l(Y-X))\times I)=\mu(\widetilde{Q}_{r}(Y)\times I).$
@@
Given that $(Y+2\ell(Y-X),s)\in\operatorname{supp}(\mu)$ for all $\ell\in\mathbb{Z}$ (see Lemma 5.3), the statement of the lemma now follows (for instance) from relabelling. ∎
@@
Given an integer $l>0$ and $(l+1)$ points $(Y_{0},s_{0}),...,(Y_{l},s_{l})\in\operatorname{supp}(\mu)$, for any $a_{1},...,a_{l}\in 2^{l}\mathbb{Z}$ the point
$\Bigl(Y_{0}+\sum_{i=1}^{l}a_{i}(Y_{i}-Y_{0}),s_{0}\Bigl)$
@@
We use induction. Without loss of generality, we may assume $(Y_{0},s_{0})=(0,0)$. Notice that the base case $l=1$ is a direct consequence of Lemma 5.3.
Assume the Proposition is known for $l=j$. We will show it holds for $l=j+1$. Choose $a_{1},...,a_{j+1}\in 2^{j+1}\mathbb{Z}$. By our induction hypothesis, the point $\bigl(\frac{1}{2}\sum_{i=1}^{j}a_{i}Y_{i},0\bigl)$ belongs to $\operatorname{supp}(\mu)$. Moreover, by Lemma 5.3, $(-a_{j+1}Y_{j+1},0)\in\operatorname{supp}(\mu)$. So, an application of Lemma 5.3 yields that $\bigl(\sum_{i=1}^{j+1}a_{i}Y_{i},0\bigl)\in\operatorname{supp}(\mu),$ as required. ∎
@@
Without loss of generality, suppose that $(X_{0},t_{0})=(0,0).$ We first claim that, because $(Y,0)\in\operatorname{supp}(\mu)$, we have $(V-Y)\times\{0\}\subseteq\operatorname{supp}(\mu).$ To see this, fix $(v-Y,0)\in(V-Y)\times\{0\}$. Then, since $(Y,0)=(v/2+(Y-v/2,0)\in\operatorname{supp}(\mu)$ and $(v/2,0)\in\operatorname{supp}(\mu)$, Lemma 5.4 implies that
@@
Finally, given $(v+Y,0)\in(V+Y)\times\{0\}$, we observe that $(0,0)$ and (vY)(-v-Y) belong to $\operatorname{supp}(\mu)$. Therefore, using Lemma 5.4 once again,
@@
Suppose that $\mu$ is a $(k+2)$-ADR regular vampiric measure, and $L$ is an $m$-dimensional spatial plane such that $\operatorname{supp}(\mu)_{t}\cap\widetilde{Q}(\mathbf{X},3r)=L$ with $X\in L$, $\mathbf{X}=(X,t)$. Then
$\mu|_{Q(\mathbf{X},r)}=\bigl(\mathcal{H}^{m}|_{L}\times\nu\bigl)|_{Q(\mathbf{X},r)}$
for a ($(k+2-\ell)$-ADR) measure $\nu$ on $\mathbb{R}$.
@@
If there exists $t\in T$ and $Y\in\widetilde{Q}(0,r)\cap\operatorname{supp}(\mu)_{t}\setminus L$, then Lemma 5.3 shows that $2Y\in\operatorname{supp}(\mu)_{0}$. Since $2Y\in\widetilde{Q}(\mathbf{0},3r)$, this contradicts our assumption that $\operatorname{supp}(\mu)_{0}\cap\widetilde{Q}(0,3r)=L\cap\widetilde{Q}(0,3r)$. Therefore
(5.7) $\operatorname{supp}(\mu)_{t}\cap\widetilde{Q}(0,r)\supset L\cap\widetilde{Q}(0,r)\text{ for every }t\in T.$
On the other hand, Lemma 5.3 ensures that since $L\cap\widetilde{Q}(0,3r)\subset(\operatorname{supp}(\mu))_{0}\cap\widetilde{Q}(0,3r)$, then in fact
@@
For $t\in T$, there exists $X\in L$ such that $(X,t)\in\operatorname{supp}(\mu)$ (recall (5.7)). Using (5.8), we infer from Lemma 5.3 that the point $(2Y-X,t)\in\operatorname{supp}(\mu)$ for every $Y\in L(\subset\operatorname{supp}(\mu)_{0})$. Since $L$ is a linear subspace this ensures that $L\subset\operatorname{supp}(\mu)_{t}$, and therefore $\operatorname{supp}(\mu)_{t}\cap\widetilde{Q}(0,r)=L\cap\widetilde{Q}(0,r)$ for every $t\in T$.
Now, for an interval $I\subset(-r,r)$, and a spatial cube $\widetilde{Q}$ centered on $L\cap\widetilde{Q}(0,r)$ and contained in $\widetilde{Q}(0,r)$, the quantity $\mu(\widetilde{Q}\times I)$ does not depend on the location of $\widetilde{Q}$ (recall Lemma 5.5). Therefore, the measure $\sigma(E)=\mu(E\times I)$ is a constant multiple of the $m$-dimensional Hausdorff measure on $L\cap\widetilde{Q}(0,r)$. So $\mu(E\times I)=c_{m}\mathcal{H}^{m}(E)\mu(\widetilde{Q}_{1}(0)\times I)$.
Therefore, well-known uniqueness results for Borel measures ensure that, when restricted to $Q(\mathbf{0},r)$, $\mu=\mathcal{H}^{m}\times\nu$. The lemma is proved.∎
@@
By way of contradiction, assume $m<k$. Because $\mu$ is $(k+2)$-ADR, every ball $B_{r}(\mathbf{0})$ contains a point $(X,t)\in\operatorname{supp}(\mu)$ such that $X\notin L_{l}$. So, we can produce a sequence $\{(X_{j},t_{j})\}_{j\geq 1}\subseteq\operatorname{supp}(\mu)$ such that $X_{j}\notin L$ for each $l$, and so that $(X_{j},t_{j})$ converges to $\mathbf{0}$. An application of Lemma 5.8 then shows that $L$ cannot be of maximal dimension, a contradiction.
@@
    There is a $m$-dimensional linear subspace $L\subset\operatorname{supp}(\mu)_{0}$ with $m=\dim L\in\{k,k+1,k+2\},$ and
@@
    a (possibly empty) set of linearly independent set of vectors $e_{1},\dots,e_{q}$ orthogonal to $L$, where $m+q\leq k+2$, and
@@
      $\operatorname{supp}(\mu)_{0}\cap\widetilde{Q}(\mathbf{0},|e_{1}|/\sqrt{n})=L\cap\widetilde{Q}(\mathbf{0},|e_{1}|/\sqrt{n}),$ and
@@
Suppose $\L\operatorname{supp}(\mu)_{0}\backslash$ $\neq\varnothing. L$emma 5.8 ensures that, if $e$ is a vector orthogonal to $L$ with $e\subset\operatorname{supp}(\mu)_{0}$, then $|e|$ is bounded from below. Select $e_{1}\in\operatorname{supp}(\mu)_{0}$ to be a vector orthogonal to $L$ of minimal length, and then, given select $e_{1},e_{2},\dots,e_{\ell}$, select (if possible) a vector $e_{\ell+1}\in\operatorname{supp}(\mu)_{0}$ of minimal length that is orthogonal $L$ and linearly independent of $e_{1},e_{2},\dots,e_{\ell}$. This process must terminate with a finite set of vectors $e_{1},\dots,e_{q}$. By construction we have that property (b) holds.
@@
It remains to check that $q\leq k+2-m$. Fix $K>1$ large. Consider $A:=\{(a_{1},...,a_{q}:a_{j}\in 2^{q}\mathbb{Z}\cap[-K,K])\}$, and set $\overline{e}=(e_{1},...,e_{q})$. Notice that $A\approx K^{q}$.
For a small parameter $\varkappa$ to be specified (chosen depending on $e_{1},\dots,e_{q}$), let $\{Q_{j}\}_{1}^{M}$ be a collection of cubes of side-length $\varkappa$ which are pairwise disjoint and such that the union of the closures of the spatial cubes $\widetilde{Q}_{j}$ covers $L_{l}\cap\widetilde{B}_{K}(0)$, and so that each cube $\widetilde{Q}_{j}$ is centered on $L\cap\widetilde{B}_{K}(0)$. We may ensure these cubes are contained in $\widetilde{B}_{2K}(0)$, in which case $M\approx K^{m}\varkappa^{-m}$.
@@
$\mu(Q_{3K}(\mathbf{0}))\gtrsim K^{m+q}\varkappa^{-k}\varkappa^{k+2}$
@@
For every $\lambda>0$ there exists $M_{1}>0$ such that the following holds: Suppose that $\mu$ is a $(k+2)$-ADR vampiric measure such that there is a $m$-dimensional linear subspace $L\subset\operatorname{supp}(\mu)_{0}$ with $m=\dim L\in\{k,k+1\},$ and a vector $e$ orthogonal to $L$ with $|e|\leq 3\sqrt{n}$ and
(5.9) $\begin{split}\widetilde{Q}(\mathbf{0},5M_{1})\cap\operatorname{span}(L,e)&\supset\widetilde{Q}(\mathbf{0},5M_{1})\cap\operatorname{supp}(\mu)_{0}\\ &\supset\widetilde{Q}(\mathbf{0},5M_{1})\cap\Bigl\{L+ae\text{ for any }a\in 4\mathbb{Z}\Bigl\}.\end{split}$
Then
$\alpha_{\mu,n+1}(B(\mathbf{0},M_{1}))<\lambda.$
@@
Suppose not! Then there exists $\lambda>0$ such that for all $M_{\ell}$ there exists $\widetilde{\mu}_{\ell}$ a $(k+2)$-ADR vampiric measure, an $m_{\ell}$-dimensional subspace $L_{\ell}$ and an orthogonal vector $e_{\ell}$ with $|e_{\ell}|\leq 3\sqrt{n}$, but $\alpha_{\widetilde{\mu},k+1}(B(\mathbf{0},M_{k}))\geq\lambda.$
Since $m_{\ell}\in\{k,k+1\}$ ($k+2$ is not possible) we may pass to a subsequence if necessary and assume that $m_{\ell}$ is constant (equal to $m$). Additionally, by passing to a further subsequence we may assume that $e_{\ell}/|e_{\ell}|\to e\in\mathbb{R}^{n}$ as $\ell\to 0$.
Now consider the measures $\mu_{\ell}=\frac{\widetilde{\mu}_{\ell}(M_{\ell}\,\cdot\,)}{M_{\ell}^{k+2}}$. These measures are $(k+2)$-ADR and vampiric, so by passing to another subsequence we may assume converge weakly to a $(k+2)$-ADR measure (Lemma 2.1) $\mu$ that is vampiric. By weak continuity of the $\alpha$-numbers, we have that $\alpha_{\mu,m+1}(B(\mathbf{0},1))\geq\lambda$. In $\widetilde{Q}(\mathbf{0},5)$, we obtain from (5.9) that $\operatorname{supp}(\mu_{\ell})_{0}$ is contained $\text{span}(L_{\ell},e_{\ell})$, and contains $\bigl\{L_{\ell}+\frac{a}{M_{\ell}}e_{\ell}\,:\,a\in 4\mathbb{Z}\bigl\}$. From this, along with the Ahlfors regularity of $\mu_{\ell}$ and $\mu$, we conclude that $\operatorname{supp}(\mu)_{0}\cap\widetilde{Q}(\mathbf{0},5)=P\cap\widetilde{Q}(\mathbf{0},5)$, where $P=\text{span}(L,e)$. But now from Lemma 5.9, we have that, in $Q(\mathbf{0},1)$,
@@
but then $\alpha_{\mu,m+1}(B(\mathbf{0},1))=0$, which is absurd. ∎

### v2

@@ 5. Vampiric measures
In light of Theorem 4.3, the task of proving Theorems 3.1 and 3.2 have been reduced to proving Carleson packing conditions on the collection of cubes $E\in\mathbb{D}(\mu)$ with noticeable $\alpha_{\mu,k}(E)$ coefficient. These theorems are provided by studying vampiric measures. The majority of the analysis takes place in the parabolic setting. The situation with the Heisenberg group will then be reduced to the parabolic case by appealing to the analysis in Orponen [51]. It is important to note that we are not only interested in the support of a vampiric measure but also its distribution.
We will call a Borel measure $\mu$ on $\mathbb{R}^{n+1}$ vampiric [^9] These measures correspond to the local symmetry condition introduced by David and Semmes [20], and so it might be natural to refer to them as symmetric measures. However, Mattila has already introduced a (wider) class of measures in the context of understanding the geometric consequences of the existence of principal value singular integrals called symmetric measures, so we decided not to use this terminology. if
@@ 5.1. The parabolic case
For every $\lambda>0$ there exist $M_{1}>0$ and $M_{2}$ such that the following holds. Suppose that a vampiric measure $\mu$ is $(k+2)$-ADR, $\mathbf{0}\in\operatorname{supp}(\mu)$ and
(5.2) $\alpha_{k,\mu}(\mathbf{0},r)>0.$
@@
(5.3) $\alpha_{k+1,\mu}(\mathbf{0},M_{1}r)<\lambda$
@@
(5.4) $\alpha_{k+2,\mu}(\mathbf{0},M_{2}r)<\lambda.$
This proposition holds with $\mathbf{0}$ replaced by any point in the support of the measure: all of the properties the measure in the statement are translation invariant. This translation invariance property will be used routinely in this section.
The proof of this proposition will require a fair amount of preparatory work. Recall the definitions of cylinders $C_{r}(\mathbf{X})$ and spatial cubes $\widetilde{Q}_{r}(X)$ from Section 2.1.
@@
Suppose $(X,t),(Y,s)\in\operatorname{supp}(\mu)$. With Z=YXZ=Y-X, the points $(X+2\ell Z,t)$, $\ell\in\mathbb{Z}$, are contained in $\operatorname{supp}(\mu)$. Moreover, for all $0<r<\infty$ we have
$\mu(C_{r}(X+2\ell(Y-X),t))=\mu(C_{r}(X,t)).$
@@
With a view to arguing by induction, consider the following statement: Suppose that $j\in\{0,1,2,\dots\,\}$ and that $(X+2\ell(Y-X),t)\in\operatorname{supp}(\mu)$ and $(Y+2\ell(Y-X),s)\in\operatorname{supp}(\mu)$ for every $\ell\in\mathbb{Z}$ with $|\ell|\leq j$, and $\mu(C_{r}(X+2l(Y-X),t))=\mu(C_{r}(X,t))$ for $\ell\in\mathbb{Z}$ with $|\ell|\leq j$.
This statement holds trivially for $j=0$. Suppose now it holds for some fixed $j$, then
$C_{r}(X+2(j+1)(Y-X),t)=(Y+2j(Y-X),s)+C_{r}((Y-X),t-s),$
so equation (5.5) with $(X,t)$ replaced by $(Y+2j(Y-X),s)\in\operatorname{supp}(\mu)$, $\xi=Y-X$, and $\tau=t-s$, yields
$\mu(C_{r}(X+2(j+1)(Y-X),t))$ $=\mu((Y+2j(Y-X),s)+C_{r}((X-Y),t-s))$
$=\mu(C_{r}(X+2j(Y-X),t))>0.$
@@
$C_{r}(X+2(j+1)(Y-X),t)=(X,t)+C_{r}(2(j+1)(Y-X),0),$
so employing (5.5), we deduce that $\mu(C_{r}(X+2(j+1)(Y-X),t))=\mu(C_{r}(X-2(j+1)(Y-X),t))=\mu(C_{r}(X,t))>0.$ Interchanging the roles of $(X,t)$ and $(Y,s)$ in the above argument now yields that $(Y\pm 2(j+1)(Y-X),s)\in\operatorname{supp}(\mu)$. This completes the induction step, and proves the lemma. ∎
@@
Suppose $(X,t),(Y,t)\in\operatorname{supp}(\mu)$. With Z=YXZ=Y-X, the points $(X+\ell Z,t)$, $\ell\in\mathbb{Z}$, are contained in $\operatorname{supp}(\mu)$.
@@
Suppose $(X,t),(Y,s)\in\operatorname{supp}(\mu)$, $r>0$, and $I$ is an open interval in $\mathbb{R}$. With $\widetilde{Q}_{r}(Z):=\{Z^{\prime}\in\mathbb{R}^{n}:\|Z-Z^{\prime}\|_{\infty}<r\}$, for each $\ell\in\mathbb{Z}$ we have
$\mu(\widetilde{Q}_{r}(X+2\ell(Y-X))\times I)=\mu(\widetilde{Q}_{r}(Y)\times I).$
@@
Given that $(Y+2\ell(Y-X),s)\in\operatorname{supp}(\mu)$ for all $\ell\in\mathbb{Z}$ (see Lemma 5.3), the statement of the lemma now follows (for instance) from repetition of the induction step in the proof of Lemma 5.3. ∎
@@
Given an integer $\ell>0$ and $(\ell+1)$ points $(Y_{0},s_{0}),...,(Y_{\ell},s_{\ell})\in\operatorname{supp}(\mu)$, for any $a_{1},...,a_{\ell}\in 2^{\ell}\mathbb{Z}$ the point
$\Bigl(Y_{0}+\sum_{i=1}^{\ell}a_{i}(Y_{i}-Y_{0}),s_{0}\Bigl)$
@@
We use induction. Without loss of generality, we may assume $(Y_{0},s_{0})=(0,0)$. Notice that the base case $\ell=1$ is a direct consequence of Lemma 5.3.
Assume the Proposition is known for $\ell=j$. We will show it holds for $\ell=j+1$. Choose $a_{1},...,a_{j+1}\in 2^{j+1}\mathbb{Z}$. By our induction hypothesis, the point $\bigl(\frac{1}{2}\sum_{i=1}^{j}a_{i}Y_{i},0\bigl)$ belongs to $\operatorname{supp}(\mu)$. Moreover, by Lemma 5.3, $(-a_{j+1}Y_{j+1},0)\in\operatorname{supp}(\mu)$. So, an application of Lemma 5.3 yields that $\bigl(\sum_{i=1}^{j+1}a_{i}Y_{i},0\bigl)\in\operatorname{supp}(\mu),$ as required. ∎
@@
Without loss of generality, suppose that $(X_{0},t_{0})=(0,0).$ We first claim that, because $(Y,0)\in\operatorname{supp}(\mu)$, we have $(V-Y)\times\{0\}\subseteq\operatorname{supp}(\mu).$ To see this, fix $(v-Y,0)\in(V-Y)\times\{0\}$. Then, since $(Y,0)=(v/2+(Y-v/2),0)\in\operatorname{supp}(\mu)$ and $(v/2,0)\in\operatorname{supp}(\mu)$, Lemma 5.4 implies that
@@
Finally, given $(v+Y,0)\in(V+Y)\times\{0\}$, we observe that $(0,0)$ and (vY,0)(-v-Y,0) belong to $\operatorname{supp}(\mu)$. Therefore, using Lemma 5.4 once again,
@@
Suppose that $\mu$ is a $(k+2)$-ADR vampiric measure, and $L$ is an $m$-dimensional spatial plane such that $\operatorname{supp}(\mu)_{t}\cap\widetilde{Q}_{3r}(X)=L$ with $X\in L$, $\mathbf{X}=(X,t)\in\operatorname{supp}(\mu)$. Then
$\mu|_{Q_{r}(\mathbf{X})}=\bigl(\mathcal{H}^{m}|_{L}\times\nu\bigl)|_{Q_{r}(\mathbf{X})}$
for a (k+2m)(k+2-m)-ADR measure $\nu$ on $\mathbb{R}$.
@@
If there exists $t\in T$ and $Y\in\widetilde{Q}_{r}(0)\cap\operatorname{supp}(\mu)_{t}\setminus L$, then Lemma 5.3 shows that $2Y\in\operatorname{supp}(\mu)_{0}$. Since $2Y\in\widetilde{Q}_{3r}(0)$, this contradicts our assumption that $\operatorname{supp}(\mu)_{0}\cap\widetilde{Q}_{3r}(0)=L\cap\widetilde{Q}_{3r}(0)$. Therefore
(5.7) $\operatorname{supp}(\mu)_{t}\cap\widetilde{Q}_{r}(0)\subset L\cap\widetilde{Q}_{r}(0)\text{ for every }t\in T.$
On the other hand, Lemma 5.3 ensures that since $L\cap\widetilde{Q}_{3r}(0)\subset\operatorname{supp}(\mu)_{0}\cap\widetilde{Q}_{3r}(0)$, and hence
@@
For $t\in T$, there exists $X\in L$ such that $(X,t)\in\operatorname{supp}(\mu)$ (recall (5.7)). Using (5.8), we infer from Lemma 5.3 that the point $(2Y-X,t)\in\operatorname{supp}(\mu)$ for every $Y\in L(\subset\operatorname{supp}(\mu)_{0})$. Since $L$ is a linear subspace this ensures that $L\subset\operatorname{supp}(\mu)_{t}$, and therefore $\operatorname{supp}(\mu)_{t}\cap\widetilde{Q}_{r}(0)=L\cap\widetilde{Q}_{r}(0)$ for every $t\in T$.
Now, for an interval $I\subset(-r^{2},r^{2})$, and a spatial cube $\widetilde{Q}$ centered on $L\cap\widetilde{Q}_{r}(\mathbf{0})$ and contained in $\widetilde{Q}_{r}(\mathbf{0})$, the quantity $\mu(\widetilde{Q}\times I)$ does not depend on the location of $\widetilde{Q}$ (recall Lemma 5.5). It follows that for Borel measurable $E\subset\widetilde{Q}_{r}(\mathbf{0})\cap L$, $\mu(E\times I)=c_{m}\mathcal{H}^{m}(E)\mu(\widetilde{Q}_{1}(\mathbf{0})\times I)$ from standard measure theoretic considerations [^10] For instance by first obtaining the equality for a (Euclidean) dyadic cube, then for open sets, and finally for Borel sets..
[…]
@@
By way of contradiction, assume $m<k$. Because $\mu$ is $(k+2)$-ADR, every ball $B_{r}(\mathbf{0})$ contains a point $(X,t)\in\operatorname{supp}(\mu)$ such that $X\notin L$. So, we can produce a sequence $\{(X_{j},t_{j})\}_{j\geq 1}\subseteq\operatorname{supp}(\mu)$ such that $X_{j}\notin L$ for each $j$, and so that $(X_{j},t_{j})$ converges to $\mathbf{0}$. An application of Lemma 5.8 then shows that $L$ cannot be of maximal dimension, a contradiction.
@@
    There is a $m$-dimensional linear subspace $L\subset\operatorname{supp}(\mu)_{0}$ with $m=\dim L\in\{k,k+1,k+2\},$
@@
    a (possibly empty) set of linearly independent vectors $e_{1},\dots,e_{q}$ orthogonal to $L$, where $m+q\leq k+2$, and
@@
      $\operatorname{supp}(\mu)_{0}\cap\widetilde{Q}_{|e_{1}|/\sqrt{n}}(0)=L\cap\widetilde{Q}_{|e_{1}|/\sqrt{n}}(0),$ and
@@
Suppose $\L\operatorname{supp}(\mu)_{0}\backslash$ $\neq\varnothing. L$emma 5.8 ensures that, if $e$ is a vector orthogonal to $L$ with $e\in\operatorname{supp}(\mu)_{0}$, then $|e|$ is bounded away from 0. Select $e_{1}\in\operatorname{supp}(\mu)_{0}$ to be a vector orthogonal to $L$ of minimal length, and then, given $e_{1},e_{2},\dots,e_{\ell}$, select (if possible) a vector $e_{\ell+1}\in\operatorname{supp}(\mu)_{0}$ of minimal length that is orthogonal $L$ and linearly independent of $e_{1},e_{2},\dots,e_{\ell}$. This process must terminate with a finite set of vectors $e_{1},\dots,e_{q}$. By construction we have that property (b) holds.
@@
It remains to check that $q\leq k+2-m$. Fix $K>1$ large. Consider $A:=\{(a_{1},...,a_{q}):a_{j}\in 2^{q}\mathbb{Z}\cap[-K,K])\}$, and set $\overline{e}=(e_{1},...,e_{q})$. Notice that, because $q\leq n$, $A\approx K^{q}$.
For a small parameter $\varkappa$ to be specified (chosen depending on $e_{1},\dots,e_{q}$), let $\{\tilde{Q}_{j}\}_{1}^{M}$ be a collection of cubes the spatial component $\mathbb{R}^{n}$ of side-length $\varkappa$ which are pairwise disjoint and such that the union of the closures of the spatial cubes $\widetilde{Q}_{j}$ covers $L\cap\widetilde{Q}_{K}(0)$, and so that each cube $\widetilde{Q}_{j}$ is centered on $L\cap\widetilde{Q}_{K}(0)$. We may ensure these cubes are contained in $\widetilde{Q}_{2K}(0)$, in which case $M\approx K^{m}\varkappa^{-m}$.
@@
$\mu(Q_{3K}(\mathbf{0}))\gtrsim K^{m+q}\varkappa^{-m}\varkappa^{k+2}$
@@
For every $\lambda>0$ there exists $M_{1}>0$ such that the following holds for every $M\geq M_{1}$: Suppose that $\mu$ is a $(k+2)$-ADR vampiric measure such that $\mathbf{0}\in\operatorname{supp}(\mu)$, there is a $m$-dimensional linear subspace $L\subset\operatorname{supp}(\mu)_{0}$ with $m=\dim L\in\{k,k+1\},$ and a vector $e$ orthogonal to $L$ with $|e|\leq 3\sqrt{n}$, and
(5.9) $\begin{split}\widetilde{Q}_{5M}(0)\cap\operatorname{span}(L,e)&\supset\widetilde{Q}_{5M}(0)\cap\operatorname{supp}(\mu)_{0}\\ &\supset\widetilde{Q}_{5M}(0)\cap\Bigl\{L+ae\text{ for any }a\in 4\mathbb{Z}\Bigl\}.\end{split}$
Then,
$\alpha_{m+1,\mu}(\mathbf{0},M)<\lambda.$
@@
Suppose not. Then there exists $\lambda>0$ and a sequence of $(k+2)$-ADR vampiric $\{\widetilde{\mu}_{\ell}\}_{\ell\in\mathbb{N}}$ measures with the following properties.
  • There exists a sequence $\{m_{\ell}\}_{\ell\in\mathbb{N}}$ with $m_{\ell}\in\{k,k+1\}$ along with a sequence of $m_{\ell}$-dimensional subspaces $\{L_{\ell}\}_{\ell\in\mathbb{N}}$ with $L_{\ell}\subset\operatorname{supp}(\widetilde{\mu}_{\ell})_{0}$, and a sequence of vectors $\{e_{\ell}\}_{\ell\in\mathbb{N}}$ satisfying $e_{\ell}\perp L_{\ell}$ and $|e_{\ell}|\leq 3\sqrt{n}$.
  • There exists an increasing sequence $\{M_{\ell}\}_{\ell\in\mathbb{N}}\subset\mathbb{R}$ such that $\alpha_{\widetilde{\mu}_{\ell},m_{\ell}+1}(\mathbf{0},M_{\ell})\geq\lambda$.
  • For each $\ell$, display 5.9 holds with $\widetilde{\mu}_{\ell}$, $M_{\ell}$, $e_{\ell}$, and $L_{\ell}$ in place of $\mu$, $M_{1}$, $e$ and $L$, respectively.
Since $m_{\ell}\in\{k,k+1\}$ we may pass to a subsequence if necessary and assume that $m_{\ell}$ is constant (equal to $m$). Additionally, by passing to a further subsequence we may assume that $e_{\ell}/|e_{\ell}|\to e\in\mathbb{R}^{n}$ as $\ell\to 0$, and that the planes $L_{\ell}$ converge to some plane $L$ in the local Hausdorff metric.
Now consider the measures $\mu_{\ell}=\frac{\widetilde{\mu}_{\ell}(M_{\ell}\,\cdot\,)}{M_{\ell}^{k+2}}$. These measures are $(k+2)$-ADR and vampiric, so by passing to another subsequence we may assume converge weakly to a $(k+2)$-ADR measure (Lemma 2.1) $\mu$ that is vampiric. By weak continuity of the $\alpha$-numbers, we have that $\alpha_{\mu,m+1}(B_{1}(\mathbf{0}))\geq\lambda$. In $\widetilde{Q}_{5}(\mathbf{0})$, we obtain from (5.9) that $\operatorname{supp}(\mu_{\ell})_{0}$ is contained $\text{span}(L_{\ell},e_{\ell})$, and contains $\bigl\{L_{\ell}+\frac{a}{M_{\ell}}e_{\ell}\,:\,a\in 4\mathbb{Z}\bigl\}$. From this, along with the Ahlfors regularity of $\mu_{\ell}$ and $\mu$, we conclude that $\operatorname{supp}(\mu)_{0}\cap\widetilde{Q}_{5}(\mathbf{0})=P\cap\widetilde{Q}_{5}(\mathbf{0})$, where $P=\text{span}(L,e)$. Notice that this is where we use the assumption $|e_{\ell}|\leq 3\sqrt{n}$. But now from Lemma 5.9, we have that, in $Q_{1}(\mathbf{0})$,
@@
but then $\alpha_{m+1,\mu}(\mathbf{0},M_{1})=0$, which is absurd. ∎